A 20- ftft-long column is made of aluminum alloy 2014-T6.

If it is pinned at its top and bottom, and a compressive load PP is applied at point AA, determine the maximum allowable magnitude of PP using the equations σallow=[54000(KL/r)2]σallow=[54000(KL/r)2] and P/A(σa)allow+Mc/Ar2(σb)allowP/A(σa)allow+Mc/Ar2(σb)allow, with (σb)allow(σb)allow = 22 ksiksi .

Express your answer to three significant figures and include appropriate units.Problem 13.118 11 of 21 A 20-ft-long column is made of aluminum alloy 2014-T6. A 5.25 in. 0.5 in. X x 10 in. 0.5 in. 0.5 in.
Answer #1

Calculate the sectional area of the column. A = (bh), -(bh). A =(10)(10+0.5+0.5)-(10)(10-0.5) A =(10x11)-(10x9.5) A = 15 in.

Calculate the slenderness ratio of the of the aluminum alloy 2014-T6 about y-y axis. KL I KL 1(20)(12) 2.358 T y KL =101.76 r